OpenStudy (anonymous):

Find a cubic function, in the form f(x)=ax^3+bx^2+cx+d, that has a local maximum value of 4 at -3 and a local minimum value of 0 at 2. f (x) = ax3 + bx2 + cx + d f′(x)=3ax2+2bx+c f(−3)=4=>−27a+9b−3c+d f(2)=0=>a+b+c+d f′(−3)=0=>27a−6b+c f′(2)=0=>12a+4b+c

OpenStudy (anonymous):

(−27a+9b−3c+d=4−a+b+c+d=0)=−28a+8b−4c=4 (27a−6b+c=0−12a+4b+c=0)=12a−10b=0 (−28a+8b−4c=4+4(12a+4b+c)=0)=(−28a+8b−4c=4+48a+16b+4c=0) =20a+24b=4 I need to isolate a and b in order to get d, then c